OpenStudy (anonymous):

Factor x^3-x^2-5x+2 completely.

OpenStudy (jtvatsim):

is this an algebra class or precalc?

OpenStudy (jtvatsim):

if it's precalc, then we have possible rational roots of +2 or -2. Testing these roots, we see that -2 works nicely, (-2)^3 -(-2)^2-5(-2)+2 = -8-4+10+2 = 0. So, (x - (-2)) = (x + 2) is one factor. Then, using synthetic division gives: -2| 1 -1 -5 2 |____-2__6__-2__ 1 -3 1 0 Using the bottom line, we have coefficients of 1, -3 and 1 so x^2 - 3x + 1 Thus, the answer is (x + 2)(x^2 -3x +1).

OpenStudy (anonymous):

its college algebra trig

OpenStudy (jtvatsim):

otherwise we have to use trial and error. :(

OpenStudy (anonymous):

we havent learned synthetic division yet, it should be simpler i guess

OpenStudy (raden):

x^3-x^2-5x+2 (x^2.....1)(x......2) do trial error untill completing its factor

OpenStudy (jtvatsim):

in general though a good trick when you have a cubic equation is to try a factor of (x + factors of the last number) or (x - factors of the last number). In this case the last number was 2, which has factors of 2 and 1, so I tried (x + 2) and (x - 2), but you could have also tried (x + 1), (x-1) they just don't work.

OpenStudy (anonymous):

ah alright, i think thats right because we were listing factors in class. thanks.

OpenStudy (whpalmer4):

When you have *any* polynomial with a leading coefficient of 1 and integer coefficients, you should try factoring out \(( x\pm n\)) where n is the various factors of the final term. You can see why this is: Assume we have a polynomial with roots a, b. \[(x-a)(x-b) = x^2 -ax -bx + ab\]That last term is generated only by the multiplication of the roots. Let's add another root, c. \[(x-c)(x^2 -ax -bx + ab) = x^3 -ax^2 -bx^2 + abx -cx^2 +acx + bcx -abc\]Again, the final term is generated only by the multiplication of the roots.

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

yeah got it thanks